OpenStudy (anonymous):

If g(x)=-x^2+6x-1 . What's g(3) ?

OpenStudy (anonymous):

-3^2+6*3-1=8

OpenStudy (anonymous):

Thanks .

OpenStudy (shaik0124):

8

OpenStudy (anonymous):

Show Work ?

OpenStudy (shaik0124):

K WAIT

OpenStudy (anonymous):

I'm Getting A Different Answer .

OpenStudy (shaik0124):

-3^2+6*3-1=-9+18-1=8

OpenStudy (shaik0124):

WHAT DO U SAY TIERRAM

OpenStudy (anonymous):

I Was Multiplying The -3^2 Wrong . But I Have It Now . Thanks .

OpenStudy (shaik0124):

TIERRAM HOW CAN WE GET ACHIEVEMENTS